Grade 11Physical Chemistry

Density of 2.05Msolution of acetic acid in water is 1.02g/ml. find molality

kartik
9 Years agoGrade 11

1 Answer

Amit Kumar
9 Years ago
D=m/vLet mass = mAnd volume= vTherefore 1.02 = m/vM=1.02VNOW MOLARITY=MOLE/ VOLUMESO 2.05= MOLE/ VOLUMEMOLE= 2.05VMOLALITY= MOLE OF SOLUTE / MASSOF SOLVENTMOLALITY= 2.05V/1.02V= 2